Patriot Corporation acquired 80 percent ownership of Seahawk Corporation on January 1, 20X8, for $200,000. At that date, Seahawk reported common stock outstanding of $75,000 and retained earnings of $150,000. The fair value of the noncontrolling interest was $50,000. The differential is assigned to equipment, which had a fair value $25,000 greater than book value and a remaining economic life of five years at the date of the business combination. Seahawk reported net income of $40,000 and paid dividends of $20,000 in 20X8.

1) Provide the journal entries recorded by Patriot during 20X8 on its books if it accounts for its investment in Seahawk using the equity method.
2) Give the consolidating entries needed at December 31, 20X8, to prepare consolidated financial statements.
Answer #1
Ques 1
No. Particulars Debit Credit
1 Investment in Seahwak corporation stock 200000
cash 200000
2 Cash 16000
   Investment in Seahwak corporation stock 16000
(dividends 20000*80%)
3 Investment in Seahwak corporation stock 32000
   income from subsidiary 32000
(40000*80%)
4 Income from subsidiary 4000
   Investment in Seahwak corporation stock 4000
(excess amortization 25000/5 years *80%)
Ques 2
No. Particulars Debit Credit
1 Common stock 75000
Retained earnigns 150000
income from seahawk corp. 32000
NCI in NI of seahwak corp 8000
dividends declared 20000
investment in Seahawk corp 196000
NCI in NA of seahawk corp. 49000
(Basic consolidation entry)
2 Depreciation expense 5000
   income from seahawk corp 4000
   NCI in NI of Seahawk Corp 1000
(Amortised excess value of reclassification)
3 Equipment 25000
   Accumulated dep 5000
investment in seahawk corp 16000
NCI in NA of seahawk corp 4000
(excess value reclassification entry)

Workings:

NCI Patriot = Common stock Retained earnings
Beginning Book vaue 45000 180000 = 75000 150000
Net income 8000 32000 = 40000
Dividends -4000 -16000 = -20000
Ending BV 49000 196000 = 75000 170000
NCI Patriot = Equipment Accumulated dep
Beginning Book vaue 5000 20000 = 25000
Net income -1000 -4000 = -5000
Ending BV 4000 16000 = 25000 -5000